### v1

@@ Abstract
This is the final derivation we were after in this section. We have derived the Polchinski’s ERG flow equation for the single Bosonic and spinless Fermionic fields. We chose spinless Fermionic fields since their computations are much easier to deal with here, and the results can be generalized to Fermionic fields with spins.
3 Bosonic Field Local Potential Approximation.
The Polchinski’s ERG flow equation derived in the previous section in equation (26) is an integro-differential equation, in general this type of differential equation is challenging to solve, especially at non-perturbative level. In Renormalization Group one can handle such type of differential equations by making use of perturbative expansion either in powers of small coupling parameter epsilon ($\epsilon$) expansion or $\frac{1}{N}$ large $N$ parameter expansion. At non-perturbative level and in the absence of expansion parameters, one way to tackle this differential equation is to do some approximation or truncation by reducing the number of degrees of freedom. Performing derivative or momentum expansion of the effective action is one possible approximation that can be undertaken, therefore we will take the effective action $S_{int}$ from equation (21) with only Bosonic fields,
$S[\phi,\Lambda]=\sum_{m}\int\prod_{i=1}^{m}\frac{d^{d}k_{i}}{(2\pi)^{d}}\mathcal{L}_{m,0}(k_{1},\cdots,k_{m},\Lambda)\phi(k_{1})\cdots\phi(k_{m})(2\pi)^{d}\delta^{d}(k_{1}+\cdots+k_{m}),$ (27)
@@
$\mathcal{L}_{m,0}(k_{1},\cdots,k_{m},\Lambda)=\mathcal{L}_{m,0}(0,\cdots,0,\Lambda)+\frac{1}{m(m-1)}(k_{1}^{2}+\cdots+k_{m}^{2})\mathcal{L}_{m-2,0}(0,\cdots,0,\Lambda)+\mathcal{O}(k^{4})+\cdots.$ (28)
@@
where $V(\phi,\Lambda)=\mathcal{L}_{m,0}(0,\cdots,0,\Lambda)\phi(k_{1})\cdots\phi(k_{m})(2\pi)^{d}\delta^{d}(k_{1}+\cdots+k_{m})$ is a local term which contains no derivatives, this term can be viewed as a local potential term, while the $W(\phi,\Lambda)$ term and other higher order expansion terms are non-local, these terms have derivatives acting on them in position space. To project out only the local potential term from the action $S[\phi,\Lambda]$ we can make use of the Hasenfratz and Hazenfratz projection operator $P(z)$ [5]. When this projection operator is applied on the action $S[\phi,\Lambda]$ it only selects the local potential term as follows,
@@
$\partial_{t}V(z,t)=$ $\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}V^{\prime\prime}(z,t)-G_{B}^{\prime}(0)V^{\prime 2}(z,t)-d_{\phi}zV^{\prime}(z,t)+dV(z,t),$ (34)
@@
Let $U(z,t)=e^{-V(z,t)}$, then $V(z,t)=-\ln(U(z,t))$, and the differential equation above will now be written as follows
@@
$=$ $V(z,t)+g_{b}\bigg{(}\frac{1}{\lambda t}f^{(-1)}(z)+\lambda_{k}^{(0)}f_{k}^{(0)}(z)+t\lambda_{k}^{(1)}f_{k}^{(1)}(z)+\frac{t^{2}}{2!}\lambda_{k}^{(2)}f_{k}^{(2)}(z)+\cdots+\frac{t^{n}}{n!}\lambda_{k}^{(n)}f_{k}^{(n)}(z)\bigg{)},$
@@
The results obtained here show that perturbing the interaction/vertex function in the vicinity of a logarithmic fixed point where we first linearize around the homogeneous solution as shown in equation (44), plugging this interaction function into the full rescaled Polchinski’s ERG flow equation as shown in equation (45), the solution leads to an infinite set tower of Hermite differential equation in expansion parameter $t$. The solution to these Hermite differential equations is a Hermite polynomial, but the Hermite eigenvalues for each Hermite differential equation are different at each order in expansion parameter $t$ as demonstrated in equations (51), (54) and (56).
@@
The non-negative eigenvalues are, $\lambda_{0}^{(1)}=\frac{19}{4}$, $\lambda_{1}^{(1)}=\frac{13}{4}$, $\lambda_{2}^{(1)}=\frac{5}{4}$ and their corresponding eigenfunctions respectively are the same as in (64),
@@
The first two negative eigenvalues are $\lambda_{3}^{(1)}=-\frac{1}{4}$ and $\lambda_{4}^{(1)}=-\frac{5}{4}$. At this order the critical exponents are as follows,
@@
The process should now be straight forward pertaining to how the critical exponents can be calculated at each order in expansion parameter $t$, therefore we will stop the computations at this order. Due to the infinite set of Hermite differentials equations obtained earlier in (46) where they had the same Hermite polynomial as a solution but with different eigenvalues at each order in expansion parameter $t$, this also naturally forces the critical exponent to have different values at each order in expansion parameter $t$.
4 Fermion fields homogeneous and hermite differential equations
@@
$S[\psi,\bar{\psi},\Lambda]=\sum_{n=1}\int\prod_{j=1}^{n}\frac{d^{d}p_{j}d^{d}q_{j}}{(2\pi)^{d}}\mathcal{L}_{0,n}(p_{1},\cdots,p_{n};q_{1},\cdots,q_{n},\Lambda)\bar{\psi}(p_{1})\psi(q_{1})\cdots\bar{\psi}(p_{n})\psi(q_{n})\delta^{d}(\sum_{i}^{n}p_{i}+\sum_{i}^{n}q_{i}),$ (72)
@@
where $\Delta_{\psi\bar{\psi}}=\int_{p}\psi(p)\frac{\delta}{\delta\psi(-p)}+\int_{p}\bar{\psi}(p)\frac{\delta}{\delta\bar{\psi}(-p)}$ and $\Delta_{\partial_{\psi\bar{\psi}}}=d+\int\frac{d^{d}p}{(2\pi)^{d}}\psi(p)p\cdot\frac{\partial}{\partial p}\frac{\delta}{\delta\psi(-p)}+\int\frac{d^{d}p}{(2\pi)^{d}}\bar{\psi}(p)p^{\cdot}\frac{\partial}{\partial p}\frac{\delta}{\delta\bar{\psi}(-p)}$. Acting with the operator $P(z,\bar{z})$ given in equation (71) on the differential equation above, so that the operator $P(z,\bar{z})$ extracts and returns only the local potential terms in $S[\psi,\bar{\psi},t]$ while at the same time the fields $\psi$ and $\bar{\psi}$ are replaced by the Grassmann variables $z$ and $\bar{z}$ respectively as shown below,
@@
$\frac{\partial^{2}f(z,\bar{z})}{\partial z\partial\bar{z}}e^{\lambda t}=e^{\lambda t}\bigg{(}(d-\lambda)f(z,\bar{z})+\frac{\partial f(z,\bar{z})}{\partial z}\frac{\partial V}{\partial\bar{z}}+\frac{\partial V}{\partial z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}-\frac{1}{2}(d-1)\bigg{(}z\frac{\partial f(z,\bar{z})}{\partial z}+\bar{z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}\bigg{)}\bigg{)},$ (82)
@@
$\frac{\partial^{2}f(z,\bar{z})}{\partial z\partial\bar{z}}=$ $(d-\lambda)f(z,\bar{z})-\frac{\partial f(z,\bar{z})}{\partial z}z+\bar{z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}-\frac{1}{2}(d-1)\bigg{(}z\frac{\partial f(z,\bar{z})}{\partial z}+\bar{z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}\bigg{)}$ (84)
$=$ $(d-\lambda)f(z,\bar{z})+\bigg{(}1-\frac{1}{2}(d-1)\bigg{)}z\frac{\partial f(z,\bar{z})}{\partial z}+\bigg{(}1-\frac{1}{2}(d-1)\bigg{)}\bar{z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}.$
@@
$\frac{\partial^{2}f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}\partial\bar{z}^{\prime}}=\frac{(d-\lambda)}{1-\frac{1}{2}(d-1)}f(z^{\prime},\bar{z}^{\prime})+z^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}}+\bar{z}^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial\bar{z}^{\prime}}.$ (86)
@@
$\frac{\partial^{2}f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}\partial\bar{z}^{\prime}}=\frac{(d-\lambda)}{\frac{1}{2}-\frac{1}{2}(d-1)}f(z^{\prime},\bar{z}^{\prime})+z^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}}+\bar{z}^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial\bar{z}^{\prime}},$ (89)
@@
$\frac{\partial^{2}f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}\partial\bar{z}^{\prime}}=\frac{(d-\lambda)}{\frac{1}{n}-\frac{1}{2}(d-1)}f(z^{\prime},\bar{z}^{\prime})+z^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}}+\bar{z}^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial\bar{z}^{\prime}},$ (92)
@@
The same function already given in equation (87) is the solution here, and the corresponding eigenvalue $\lambda^{n)}=\lambda$ is
@@
$=$ $\bigg{(}\frac{z\bar{z}}{\lambda t}+\ln(\lambda t)\bigg{)}+g_{f}\bigg{(}\frac{1}{\lambda t}f^{(-1)}(z,\bar{z})+\lambda^{(0)}f^{(0)}(z,\bar{z})+\lambda^{(1)}tf^{(1)}(z,\bar{z})\bigg{)}$
@@
$V_{P}(z,\bar{z})=$ $\big{(}1+\ln(z\bar{z})\big{)}+g_{f}\bigg{(}\frac{1}{z\bar{z}}f^{(-1)}(z,\bar{z})+\lambda^{(0)}f^{(0)}(z,\bar{z})+\frac{\lambda^{(1)}}{\lambda}z\bar{z}f^{(1)}(z,\bar{z})\bigg{)}$ (96)
$=$ $\big{(}1+\ln(z\bar{z})\big{)}+g_{f}\bigg{(}\frac{1}{z\bar{z}}f^{(-1)}(z,\bar{z})+\lambda^{(0)}\bigg{(}z\bar{z}-\frac{1}{2}\bigg{)}-\frac{1}{2}\frac{\lambda^{(1)}}{\lambda}z\bar{z}\bigg{)}.$

### v2

@@ I Introduction
This is the final derivation we were after in this section. We have derived the Polchinski’s ERG flow equation for the single Bosonic and spinless Fermionic fields. We chose spinless Fermionic fields since their computations are much easier to handle here, and the results can be generalized to Fermionic fields with spins.
III Bosonic Field Local Potential Approximation.
The Polchinski’s ERG flow equation derived in the previous section in equation (29) is an integro-differential equation, in general this type of differential equation is challenging to solve, especially at non-perturbative level. In Renormalization Group one can handle such type of differential equations by making use of perturbative expansion either in powers of small coupling parameter epsilon ($\epsilon$) expansion or $\frac{1}{N}$ large $N$ parameter expansion. At non-perturbative level and in the absence of expansion parameters, one way to tackle this differential equation is to do some approximation or truncation by reducing the number of degrees of freedom. Performing derivative or momentum expansion of the effective action is one possible approximation that can be undertaken, therefore here we will take the effective action $S_{int}$ from equation (21) with only Bosonic fields,
$S[\phi,\Lambda]=$ $\sum_{m}\int\prod_{i=1}^{m}\frac{d^{d}k_{i}}{(2\pi)^{d}}\mathcal{L}_{m,0}(k_{1},\cdots,k_{m},\Lambda)$ (30)
×$\times\phi(k_{1})\cdots\phi(k_{m})(2\pi)^{d}\delta^{d}(k_{1}+\cdots+k_{m}),$
@@
$,\displaystyle\mathcal{L}_{m,0}(k_{1},$$\cdots,k_{m},\Lambda)=\mathcal{L}_{m,0}(0,\cdots,0,\Lambda)$ (31)
$+\frac{1}{m(m-1)}(k_{1}^{2}+\cdots+k_{m}^{2})\mathcal{L}_{m-2,0}(0,\cdots,0,\Lambda)$
$+\mathcal{O}(k^{4})+\cdots.$
@@
where
$V(\phi,\Lambda)=\mathcal{L}_{m,0}(0,\cdots,0,\Lambda)\phi(k_{1})\cdots\phi(k_{m})(2\pi)^{d}\delta^{d}(k_{1}+\cdots+k_{m})$ (33)
is a local term which contains no derivatives, this term can be viewed as a local potential term, while the $W(\phi,\Lambda)$ term and other higher order expansion terms are non-local, these terms have derivatives acting on them in position space. To project out only the local potential term from the action $S[\phi,\Lambda]$ we can make use of the Hasenfratz and Hasenfratz projection operator $P(z)$ [5]. When this projection operator is applied on the action $S[\phi,\Lambda]$ it only selects the local potential term as follows,
@@
$P(z)\big{(}\partial_{t}+d_{\phi}\triangle_{\phi}+\triangle_{\partial}-d\big{)}S[\phi,t]=P(z)\bigg{(}\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S[\phi,t]}{\delta\bar{\phi}(p)}\frac{\delta S[\phi,t]}{\delta\phi(-p)}-\frac{\delta^{2}S[\phi,t]}{\delta\bar{\phi}(p)\delta\phi(-p)}\bigg{)}\bigg{)},$ (36)
@@
$\partial_{t}V(z,t)=$ $\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}V^{\prime\prime}(z,t)-G_{B}^{\prime}(0)V^{\prime 2}(z,t)$ (38)
$-d_{\phi}zV^{\prime}(z,t)+dV(z,t),$
@@
Now let $U(z,t)=e^{-V(z,t)}$, then $V(z,t)=-\ln(U(z,t))$, and the differential equation above will now be written as follows
@@
$=$ $V(z,t)+g_{b}\bigg{[}\frac{1}{\lambda t}f^{(-1)}(z)+\lambda_{k}^{(0)}f_{k}^{(0)}(z)+t\lambda_{k}^{(1)}f_{k}^{(1)}(z)+\frac{t^{2}}{2!}\lambda_{k}^{(2)}f_{k}^{(2)}(z)+\cdots+\frac{t^{n}}{n!}\lambda_{k}^{(n)}f_{k}^{(n)}(z)\bigg{]}$
@@
The results obtained here show that perturbing the interaction or vertex function in the vicinity of a logarithmic fixed point where we first linearize around the homogeneous solution as shown in equation (48), plugging this interaction function into the full rescaled Polchinski’s ERG flow equation as shown in equation (49), the solution leads to an infinite set tower of Hermite differential equation in expansion parameter $t$. The solution to these Hermite differential equations is a Hermite polynomial, but the Hermite eigenvalues for each Hermite differential equation are different at each order in expansion parameter $t$ as demonstrated in equations (55), (58) and (60).
@@
The non-negative eigenvalues are, $\lambda_{0}^{(1)}=\frac{19}{4}$, $\lambda_{1}^{(1)}=\frac{13}{4}$, $\lambda_{2}^{(1)}=\frac{5}{4}$ and their corresponding eigenfunctions respectively are the same as in equation (68),
@@
The first two negative eigenvalues are $\lambda_{3}^{(1)}=-\frac{1}{4}$ and $\lambda_{4}^{(1)}=-\frac{5}{4}$. At this order the critical exponents will be as follows,
@@
At this point the process should now be straight forward pertaining to how the critical exponents can be calculated at each order in expansion parameter $t$, therefore we will stop the computations at this order. Due to the infinite set of Hermite differentials equations obtained earlier in equation (50) where they had the same Hermite polynomial as the solution but with different eigenvalues at each order in expansion parameter $t$, this also naturally forces the critical exponents to have different values at each order in expansion parameter $t$.
IV Fermion fields homogeneous and hermite differential equations
@@
$S[\psi,\bar{\psi},\Lambda]=$ $\sum_{n=1}\int\prod_{j=1}^{n}\frac{d^{d}p_{j}d^{d}q_{j}}{(2\pi)^{d}}\mathcal{L}_{0,n}(p_{1},\cdots,p_{n};q_{1},\cdots,q_{n},\Lambda)\times\bar{\psi}(p_{1})\psi(q_{1})\cdots\bar{\psi}(p_{n})\psi(q_{n})\delta^{d}(\sum_{i}^{n}p_{i}+\sum_{i}^{n}q_{i}),$ (76)
@@
where
$\Delta_{\psi\bar{\psi}}=\int_{p}\psi(p)\frac{\delta}{\delta\psi(-p)}+\int_{p}\bar{\psi}(p)\frac{\delta}{\delta\bar{\psi}(-p)}$ (78)
$\Delta_{\partial_{\psi\bar{\psi}}}=d+\int\frac{d^{d}p}{(2\pi)^{d}}\psi(p)p\cdot\frac{\partial}{\partial p}\frac{\delta}{\delta\psi(-p)}$
$\qquad\qquad+\int\frac{d^{d}p}{(2\pi)^{d}}\bar{\psi}(p)p^{\cdot}\frac{\partial}{\partial p}\frac{\delta}{\delta\bar{\psi}(-p)}$
Acting with the operator $P(z,\bar{z})$ given in equation (75) on the differential equation above, so that the operator $P(z,\bar{z})$ extracts and returns only the local potential terms in $S[\psi,\bar{\psi},t]$ while at the same time the fields $\psi$ and $\bar{\psi}$ are replaced by the Grassmann variables $z$ and $\bar{z}$ respectively as shown below,
@@
$\frac{\partial^{2}f(z,\bar{z})}{\partial z\partial\bar{z}}e^{\lambda t}=e^{\lambda t}\bigg{[}(d-\lambda)f(z,\bar{z})+\frac{\partial f(z,\bar{z})}{\partial z}\frac{\partial V}{\partial\bar{z}}$ (87)
$+\frac{\partial V}{\partial z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}-\frac{1}{2}(d-1)\bigg{(}z\frac{\partial f(z,\bar{z})}{\partial z}+\bar{z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}\bigg{)}\bigg{]},$
@@
$\frac{\partial^{2}f(z,\bar{z})}{\partial z\partial\bar{z}}=$ $(d-\lambda)f(z,\bar{z})+\bigg{(}1-\frac{1}{2}(d-1)\bigg{)}z\frac{\partial f(z,\bar{z})}{\partial z}$ (89)
$+\bigg{(}1-\frac{1}{2}(d-1)\bigg{)}\bar{z}\frac{\partial f(z,\bar{z})}{\partial\bar{z}}.$
@@
$\frac{\partial^{2}f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}\partial\bar{z}^{\prime}}=$ $\frac{(d-\lambda)}{1-\frac{1}{2}(d-1)}f(z^{\prime},\bar{z}^{\prime})+z^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}}$ (91)
$+\bar{z}^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial\bar{z}^{\prime}}.$
@@
$\frac{\partial^{2}f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}\partial\bar{z}^{\prime}}=$ $\frac{(d-\lambda)}{\frac{1}{2}-\frac{1}{2}(d-1)}f(z^{\prime},\bar{z}^{\prime})+z^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}}$ (94)
$+\bar{z}^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial\bar{z}^{\prime}},$
@@
$\frac{\partial^{2}f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}\partial\bar{z}^{\prime}}=$ $\frac{(d-\lambda)}{\frac{1}{n}-\frac{1}{2}(d-1)}f(z^{\prime},\bar{z}^{\prime})+z^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial z^{\prime}}$ (97)
$+\bar{z}^{\prime}\frac{\partial f(z^{\prime},\bar{z}^{\prime})}{\partial\bar{z}^{\prime}},$
@@
The same function already given in equation (92) is the solution here, and the corresponding eigenvalue $\lambda^{(n)}=\lambda$ is
@@
$=$ $\bigg{(}\frac{z\bar{z}}{\lambda t}+\ln(\lambda t)\bigg{)}+g_{f}\bigg{(}\frac{1}{\lambda t}f^{(-1)}(z,\bar{z})$
$+\lambda^{(0)}f^{(0)}(z,\bar{z})+\lambda^{(1)}tf^{(1)}(z,\bar{z})\bigg{)}$
@@
$V_{P}(z,\bar{z})=$ $\big{(}1+\ln(z\bar{z})\big{)}+g_{f}\bigg{(}\frac{1}{z\bar{z}}f^{(-1)}(z,\bar{z})$ (101)
$+\lambda^{(0)}\bigg{(}z\bar{z}-\frac{1}{2}\bigg{)}-\frac{1}{2}\frac{\lambda^{(1)}}{\lambda}z\bar{z}\bigg{)}.$